5) A book claims that more hockey players are born in January through March than in October through December. The following data show the number of players selected in a draft of new players for a hockey league according to their birth month. Is there evidence to suggest that hockey players' birthdates are not uniformly distributed throughout the year? Use the level of significance

alpha equalsα=0.05

Birth Month

Observed Count

Expected Count


January-March

60



April-June

51



July-September

26



October-December

37

4) A traffic safety company publishes reports about motorcycle fatalities and helmet use. In the first accompanying data table, the distribution shows the proportion of fatalities by location of injury for motorcycle accidents. The second data table shows the location of injury and fatalities for

2084

riders not wearing a helmet. Complete parts (a) and (b) below.

Location of injury

Observed Count

Expected Count

Multiple Locations

1049


Head

866


Neck

3535


Thorax

89


Abdomen/Lumbar/Spine

45

1)

Determine the expected count for each outcome.


n=605


i

1

2

3

4


pi

0.16

0.44

0.24

0.16


What is the expected count for outcome 1 is ( ) ?

3) The first significant digit in any number must be 1, 2, 3, 4, 5, 6, 7, 8, or 9. It was discovered that first digits do not occur with equal frequency. Probabilities of occurrence to the first digit in a number are shown in the accompanying table. The probability distribution is now known as Benford's Law. For example, the following distribution represents the first digits in

206

allegedly fraudulent checks written to a bogus company by an employee attempting to embezzle funds from his employer. Complete parts (a) through (c) below.

Distribution of first digits (Benford's Law)


Digit

1

2

3

4

5


Probability

0.301

0.176

0.125

0.097

0.079


Digit

6

7

8

9



Probability

0.067

0.058

0.051

0.046



First digits in allegedly fraudulent checks

First digit

1

2

3

4

5

6

7

8

9

Frequency

36

25

28

20

23

36

15

16

7
Answer #1

(3)

$$ \begin{array}{l} \chi^{2}=\sum\left(\frac{\left(O_{i}-E_{i}\right)^{2}}{E_{i}}\right) \\ \text { whre, } E_{i}=N * \text { Probability } \mathrm{N}=238 \end{array} $$

\(\therefore \chi^{2}=58.380\)